OpenStudy (anonymous):

Write the equation in a standard form ax^2+bx+c=0 , then give the values of a,b, and c. (x+3/2)^2 = 1/4

OpenStudy (lalaly):

\[\large{x^2+3x+\frac{9}{4}}=\large{\frac{1}{4}}\]\[\large{x^2+3x+2}=\large{0}\]

OpenStudy (lalaly):

\[\large{a}=\large{1}\]

OpenStudy (lalaly):

b=3 , c=2

OpenStudy (anonymous):

x^2+3x+2=0

OpenStudy (anonymous):

how did u get this? x^2+3x+2=0

OpenStudy (lalaly):

\[(x+\frac{3}{2})(x+\frac{3}{2})= \large{x^2+\frac{3}{2}x+\frac{3}{2}x+\frac{9}{4}}\]

OpenStudy (anonymous):

ok.. how about the values?

OpenStudy (lalaly):

\[\large{x^2+\frac{3}{2}x+\frac{3}{2}x+\frac{9}{4}}=\frac{1}{4}\]

OpenStudy (lalaly):

\[\large{x^2+3x+\frac{9}{4}}=\frac{1}{4}\]

OpenStudy (lalaly):

\[\large{x^2+3x+\frac{9}{4}-\frac{1}{4}=0}\]

OpenStudy (lalaly):

\[\large{x^2+3x+\frac{8}{4}=0}\]

OpenStudy (lalaly):

\[\large{x^2+3x+2}=0\]

OpenStudy (anonymous):

ok ... i got it.. thanks